Hello. I need help with these questions please. I need asap. Thank you. If you can, please explain how you got the answer. Thanks
1. Find the slope and the y-intercept of the line. Write fractional answers in lowest terms.
4y+7x+9= 14+7x
The slope is:
The y-intercept is (0, )
2. Find the slope if it exists of the line containing the pair of points.

(-10.3, 11.8) and (-9.9,4.3)
The slope is:

3. Find the slope and the y-intercept.

f(x)= -9x-7
The slope is:
The y-intercept is (0, )

4. Find an equation of the line having the given slope and containing the given point. Express your answer in the form x=a, y=b, or y=mx+b
m = -4,(9,0)
The equation of the line is y=

5. Write an equation of the line containing the given point and parallel to the given line. Express your answer in the form y=mx+b

(-4,9);9x=4y+7
The equation of the line is y=



Answer by mananth(16946)

1. Find the slope and the y-intercept of the line. Write fractional answers in lowest terms.
4y+7x+9= 14+7x
4y=5
y=5/4
The slope is:
The y-intercept is (0, 5/4)
2. Find the slope if it exists of the line containing the pair of points.

(-10.3, 11.8) and (-9.9,4.3)
The slope is:-18.75

3. Find the slope and the y-intercept.

f(x)= -9x-7
The slope is:-9
The y-intercept is (0, -7 )

4. Find an equation of the line having the given slope and containing the given point. Express your answer in the form x=a, y=b, or y=mx+b
m = -4,(9,0)
The equation of the line is y=-4x+36

5. Write an equation of the line containing the given point and parallel to the given line. Express your answer in the form y=mx+b

(-4,9);9x=4y+7
The equation of the line is y=(9/4)x-47/4
Hope this helps you in your online homework